# Large Language Models for Cryptocurrency Transaction Analysis: A Bitcoin Case Study

### v2

@@ 5. Discussion
Figure 6. Evaluation on different models (x axis for category, y for corresponding rates (%))
Remaining chanllegens. We stress three challenges when applying LLMs to cryptocurrency transaction graphs.
  • $\bullet$
    Token limits restrict the amount of graph data that can be processed at once. This constraint impacts the analysis of large cryptocurrency transaction graphs, as LLMs often lack sufficient contextual information to deliver accurate insights. When only a limited portion of the graph is available, it becomes challenging to capture the complexity and nuances of data.
  • $\bullet$
    The selection of reference graphs, i.e., the identification of the most representative labeled samples, is another issue. Different graphs may emphasize varying features, introducing bias. The choices can significantly influence classification outcomes and their explanations. Token limits further complicate this process. They restrict the number of samples that can be used as references, particularly for raw graph data, limiting the model’s ability to generalize its findings.
  • $\bullet$
    Improving the accuracy of LLM-generated explanations is crucial yet hard. Developing rigorous methods to enhance the precision of these explanations is necessary. Enhancing the precision of these explanations could lead to deeper analytical capabilities, potentially advancing both the interpretability and effectiveness of cryptocurrency transaction analysis.
Figure 7. Token consumption in different graph formats
Factors that may affect LLM performance? The performance is influenced by both the models used and the input data.
  • $\bullet$
    Model bottleneck. Though the major update and optimization could enhance the capability of LLMs in some aspects, it may be unable to solve the inherent problem of LLMs on transaction graph understanding. For example, using the same graph feature data for LLM-based classification, compared with GPT-3.5, the effectiveness of GPT-4 and GPT-4o improves slightly. Currently, a more effective method to boost LLMs’ performance might involve using more labeled samples and enriching the feature set in graph feature-based classifications. Feature data requires fewer LLM tokens compared to processing raw graphs.
  • $\bullet$
    Graph complexity/size. The size and complexity of transaction graphs impact LLM performance. Smaller graphs yield better foundational metric performance, while larger graphs may reduce effectiveness. At the contextual interpretation level, we found that even when using the same data, the accuracy results may differ slightly, by approximately 5%. Finally, due to the application of LLM4TG, there is a slight loss of temporal information, which may affect the results using raw graphs.
  • $\bullet$
    Data representation and features. The type of data used (e.g., graph features versus raw graphs) and the availability of labeled samples affect performance. Graph feature-based data is more efficient in terms of LLM token consumption and may improve classification effectiveness with additional features and labels.
6. Conclusion
Acknowledgment
This research is supported by OpenAI Researcher Access Program.
Appendix A LLMs Responses
Level 2 Example - bc1qah. GPT-4: $\bullet$ High in-degree and out-degree of transactions; Transactions like $n_{1}$ and $n_{2}$ have high in-degrees and out-degrees. [Average - Inaccurate] $\bullet$ Significant value differences; The in- and out-value of certain transactions show significant volume and discrepancies, such as in $n_{1}$ and $n_{2}$. [Low] GPT-4o: $\bullet$ High in-degree and out-degree of addresses in Layer 2; The addresses in Layer 2 have exceptionally high in-degree and out-degree values, indicating a large number of transactions. For example, address $n_{14}$ has an in-degree of 198 and out-degree of 188. [High] $\bullet$ High in-value and out-value in transactions; The transactions in this graph involve significant amounts of value. For instance, transaction $n_{1}$ has an in-value of 77.29740945 and out-value of 77.29452845. [High]
Level 2 Example - bc1q4w. GPT-4: $\bullet$ High out-degree in transaction nodes; A noticeable feature in this graph is the significant out-degree in some transaction nodes. For instance, $n_{1}$ has an out-degree of 600. [High] $\bullet$ Value discrepancy between in-value and out-value in transactions; The transaction node $n_{1}$ shows a discrepancy between the total incoming value (in-value: 27.69691553) and the outgoing value (out-value: 27.69184153). [Average - Invalid] GPT-4o: $\bullet$ High out-degree transactions; The transaction node $n_{1}$ has an extremely high out-degree of 600, meaning it distributes funds to 600 different address nodes. [High] $\bullet$ Large volume of transactions; The transaction node $n_{1}$ also handles a substantial volume of in-value and out-value (27.69691553). [High]
@@ Appendix A LLMs Responses
Appendix B Related Work
Empirical analysis in cryptocurrency. Empirical analysis plays an important role in understanding the dynamics of the cryptocurrency ecosystem [27, 28, 29] and the behaviors of addresses/entities [14, 30]. For instance, Tovanich et al. [31] and Hou et al. [32] reveal that factors such as payout schemes and pool fees influence miners’ behaviors in Bitcoin mining pools, and then impact the overall system performance.
Cryptocurrency analysis with advanced techniques. We introduce two methods as below.
  • $\bullet$
    Complex networks use graph theory, centrality measures, and network topology analysis to examine patterns within networks. This method explores the structures of cryptocurrency transaction networks from a macro perspective, highlighting the interconnections between nodes and the network’s overall structure.
Nerurkar et al. [38] and Serena et al. [39] drew parallels between cryptocurrency systems and other complex systems and identified characteristics such as small-world property, indicating most nodes in the graph are not neighbors but most of them can be reached by every other within a few hops. Moreover, Tao et al. [40] employed an innovative random walk with a flying-back sampling method on Bitcoin transaction graphs, uncovering phenomena such as the non-rich-club effect, i.e., that high-degree nodes are not more interconnected among themselves than with lower-degree nodes. Guo et al. [16] analyzed Ethereum transaction graphs, revealing heavy-tailed property in transaction networks, i.e., the majority of nodes have a relatively low degree while a small number of nodes have a very high degree. To mitigate the issue of money laundering on blockchain networks, DenseFlow framework, proposed by Lin et al. [41], uses dense subgraphs and the maximum flow algorithm to trace laundering activities. This approach improves precision compared to existing methods on Ethereum, demonstrating the effectiveness of network analysis in combating money laundering.
  • $\bullet$
    Machine learning is used to achieve node- or graph-level classification and prediction tasks for concrete addresses/entities.
Chaudhari et al. [42] studied utilizing temporal features to detect Bitcoin address behavioral changes and identify money laundering activities. The proposed approaches based on the decision tree (DT) by Rathore et al. [43] show high accuracy rates in detecting illicit activities and phishing scams in cryptocurrencies. Wahrstatter et al. [44] also contributed by enhancing the detection of criminal activities in Bitcoin transactions using unsupervised learning. Additionally, various machine learning methods, such as random forest (RF), multilayer perceptron (MLP), and graph neural network (GNN), are applied to benchmark cryptocurrency datasets, including the Elliptic Data Set by Weber et al. [15] and datasets by Xiang et al. [13, 25]. Besides, Gai et al. [45] proposed a transformer-based anomaly detection model BlockGPT for the Ethereum network that demonstrated acceptable utility.
LLMs in graph analysis. Several studies introduced LLMs to graph analysis, using LLM as a classifier and GNN enhancement [20]. Wang et al. [11, 17] evaluated the basic capabilities of LLMs in natural language graph problem-solving. Both studies showed limitations of LLMs, particularly in solving complex graph structures and tasks. Complementing these insights, Liu and Wu [46] and Hu et al. [10] assessed the performance of LLMs in graph data analysis and prediction, compared them with specialized GNNs. Likewise, Sui et al. [47] and Jiang et al. [48] explored the effectiveness of LLMs including GPT-3.5 and GPT-4 in processing structured data, such as tables and various structured data types, introducing innovative prompting methods for performance enhancement.
Gaps in existing studies. We conclude three primary research limitations according to the abovementioned work:
  • $\bullet$
    Existing research mainly focuses on knowledge graphs and randomly generated graphs [49, 10, 46, 11]. However, how to measure the LLMs’ ability to understand and analyze real-world cryptocurrency transaction graphs is still unresolved.
  • $\bullet$
  • $\bullet$

### v3

@@ V-D Level 3 - Contextual Interpretation
Figure 6: Evaluation on different models (x axis for category, y for corresponding rates (%))
Comparison with Traditional Models. As illustrated in Fig.6, in the context of a very small number of reference samples, LLMs based on graph features generally perform significantly better than support vector machine (SVM) and MLP, yet remain inferior compared to DT, RF, CatBoost, and GNN. Among feature-based LLMs, DeepSeek delivers the strongest results, with accuracy 46.68% and F1 score 36.56% surpassing GPT-3.5 and GPT-4, though still behind tree models and GNN.
For raw-graph inputs, performance varies widely. GPT-4 only slightly outperforms support vector machine and multilayer perceptron, whereas GPT-4o achieves substantial improvements, with accuracy 50.49% and precision 63.33% comparable to tree models. This high precision is crucial for illegal address detection, as misclassifying legitimate addresses as illegal could lead to severe consequences such as account suspension or fund freezing. DeepSeek, however, lags behind GPT-4o in both accuracy and F1, though it still demonstrates relatively strong precision 48.12%, close to some tree models.
Overall, GPT-4o analyzing sampled raw graphs attains performance close to carefully engineered tree-based methods, while DeepSeek shows that feature-driven methods can also be highly competitive.
Analysis of Classification Explanation. Unlike traditional results of classification tasks, LLM-based outcomes can include detailed explanations that may be valuable for further analysis. However, our initial investigation indicates that these explanations are not always accurate. We choose two near-accurate explanations for 1Bsjsa4441BsjsaHST2Qohs8ZHxNHeZ1UfWhtxoKHEN and 124mpe555124mpePGM2vEqcHUK96wQXur9vS7Vn7Kdj from GPT-4o to show the potential of LLMs and how they utilize different types of data.
@@
These examples illustrate that for feature-based classification, adequate quantity and quality samples with labels are required to improve classification consequences since the selected features are already processed. In contrast, for graph-based classification, the quality of samples themselves seems to be also essential. The unsuitable compression of the graphs leading to much information loss may negatively affect the results. Also, the chain of thought (CoT) prompting [52] could be used for further explanations, with detailed reasoning and better classification interpretability.
Level 3 Findings. LLMs exhibit strong capabilities in contextual interpretation, achieving high top-3 accuracy rates even with limited data sets. This suggests that they can capture broad relational patterns and generate plausible explanations. However, their overall accuracy in exact classification remains moderate, trailing behind feature-engineered tree models and GNNs. This gap indicates that while LLMs are promising for exploratory analysis and hypothesis generation, further optimization is required before they can rival specialized models in precise detection tasks.
VI Discussion
VI-A Efficiency of Graph Representation and Processing
Runtime Comparison of LLM4TG and CETraS. Table IV shows that both CETraS and LLM4TG scale nearly linearly with graph size. CETraS is consistently faster, whereas LLM4TG incurs extra overhead from its enriched representation schema. Nevertheless, both methods remain efficient, with execution times within tens of seconds even for graphs exceeding 2,000 nodes.
TABLE IV: Performance of CETraS and LLM4TG
CETraS LLM4TG
Nodes Time (s) Nodes Time (s)
49 0.0173 10 0.0009
59 0.0172 171 0.0515
385 0.1440 1806 8.3211
598 2.6626 2307 12.8064
2311 13.8108 2402 48.9033
TABLE V: LLM Capability on foundational metrics across graph formats and models
GPT-4o LLaMA DeepSeek
Metrics LLM4TG GEXF GML GraphML LLM4TG GEXF GML GraphML LLM4TG GEXF GML GraphML
struct_correctness 100.00% 95.83% 95.83% 100.00% 100.00% 87.50% 95.83% 87.50% 100.00% 95.83% 91.67% 95.83%
global_in_degree 41.67% 78.26% 78.26% 75.00% 54.17% 71.43% 69.57% 76.19% 50.00% 69.57% 50.00% 60.87%
global_out_degree 62.50% 60.87% 56.52% 54.17% 50.00% 71.43% 73.91% 76.19% 58.33% 47.83% 36.36% 39.13%
global_in_value 54.17% 30.43% 52.17% 25.00% 33.33% 33.33% 21.74% 47.62% 45.83% 56.52% 45.45% 47.83%
global_out_value 25.00% 26.09% 30.43% 25.00% 45.83% 9.52% 8.70% 19.05% 37.50% 26.09% 22.73% 26.09%
global_diff_degree 37.50% 43.48% 43.48% 41.67% 33.33% 47.62% 43.48% 52.38% 29.17% 30.43% 18.18% 34.78%
global_diff_value 8.33% 8.70% 8.70% 8.33% 25.00% 0.00% 4.35% 4.76% 16.67% 8.70% 4.55% 4.35%
node_in_degree 73.75% 75.22% 75.65% 73.33% 73.33% 66.19% 60.87% 69.05% 75.42% 73.91% 67.27% 68.26%
node_out_degree 75.83% 82.17% 80.43% 79.17% 76.25% 67.14% 54.78% 70.48% 77.50% 76.96% 72.27% 70.87%
node_in_value 75.42% 70.43% 71.74% 67.92% 73.75% 69.05% 58.26% 68.10% 77.50% 69.13% 67.27% 67.83%
node_out_value 77.08% 74.78% 71.74% 71.67% 78.33% 55.71% 55.65% 58.57% 79.17% 70.00% 68.64% 65.65%
node_special_info_a 100.00% 23.08% 14.66% 11.67% 99.17% 47.17% 47.86% 44.34% 100.00% 45.69% 46.43% 49.57%
node_special_info_t 71.67% 81.42% 64.91% 79.17% 78.33% 76.92% 72.57% 72.12% 83.90% 63.16% 64.81% 68.14%
Cross-Format Evaluation of Graph Representations. We conduct the comparison on foundational metrics shown in Table V, as they provide a direct measure of how well each format preserves graph structure and attributes. Unlike higher-level tasks, which conflate representation quality with model capacity, foundational metrics isolate the contribution of the input format itself, making them particularly suitable for cross-format evaluation.
Across three LLMs, LLM4TG achieves 100% accuracy on struct_correctness and preserves special attributes almost entirely, with node_special_info_a remaining near 100% across models. In contrast, these metrics drop in alternative formats, with struct_correctness declining to about 87–95% under LLaMA and DeepSeek, and node_special_info_a falling much more sharply to roughly 23% in GEXF, 15% in GML, and only 12% in GraphML under GPT-4o, and around 45% for LLaMA and DeepSeek.
At the node level, LLM4TG also delivers consistently stronger performance, with most metrics reaching 75–80% accuracy. For example, its node_out_value remains close to 80% across all three models, clearly surpassing the corresponding results of alternative formats. In addition, LLM4TG maintains balanced accuracy on both degree- and value-based node metrics, while alternative formats exhibit larger fluctuations and often about 60%-70%, underscoring the stability of its representation across models.
Global metrics show a more mixed pattern. LLM4TG delivers stable performance across models, avoiding the sharp fluctuations of other formats, though it is not always the top performer. In GPT-4o, global_in_degree reaches about 78% in GEXF/GML versus 42% in LLM4TG, and LLaMA performs better with GEXF/GraphML for global_out_degree with values above 70%. Still, LLM4TG remains competitive, with about 63% on global_out_degree for GPT-4o and 46% on global_in_value for DeepSeek, while keeping variance comparatively low. The global_diff_value metric is weak across all formats, never exceeding 25%, suggesting the task itself is the main challenge.
Overall, LLM4TG shows a balanced performance that achieves exact structural fidelity, preserves attributes with near-perfect accuracy, and maintains consistently strong node-level results. While global metrics remain mixed, LLM4TG avoids the sharp fluctuations of alternative formats, which exhibit uneven and task-specific strengths without offering stable benefits. These findings indicate that LLM4TG provides a concise, task-aligned representation that supports stable and robust transaction graph analysis with LLMs.
Figure 7: Token consumption in different graph formats
Token Consumption in Various Graph Formats We studied the differences in token consumption among various graph representation formats under the byte pair encoding-based tokenizer cl100k_base, the default GPT-3.5 and GPT-4 tokenizer. We chose three well-defined formats, i.e., GEXF, GML, and GraphML, for their efficiency and flexibility in representing transaction graphs. These widely used formats offer diverse encoding styles while supporting complex attributes of nodes and edges.
As in Fig.7, token usages of these formats exhibit significantly steeper curves, i.e., they consume a substantially larger number of tokens with an increasing number of nodes. These graph formats have extreme syntactic noise (i.e., great redundancy exists when representing graph data), which would result in numerous unnecessary tokens for LLMs to consume. The maximum token limits for GPT-3.5, DeepSeek, and GPT-4/4o (16,385, 64,000, and 128,000) are indicated by horizontal lines, where token usages of these formats surpass the GPT-4/4o limit when the number of nodes of a single graph exceeds approximately 500 to 750. Therefore, these graph representation formats are insufficient for LLMs to analyze transaction graphs.
VI-B LLMs for Transaction Graphs: Benefits and Challenges
Potential Advantages of LLMs. We summarise four advantages observed in applying LLMs to transaction graph analysis.
  • Robust Performance under Limited Data. LLMs can deliver meaningful results even with very limited labeled samples, achieving relatively high top-3 accuracy together with interpretable outputs. This makes them particularly useful for transaction types with scarce annotations, such as money laundering or Ponzi schemes, where traditional models often struggle due to data scarcity.
  • Context-aware Interpretation. Beyond identifying statistical patterns, LLMs can capture aspects of human intent and behavioral context that are often overlooked by traditional data-driven approaches. Leveraging their extensive training on natural language and human behavior, LLMs provide more nuanced explanations of suspicious transactions. Although these interpretations are not always fully accurate, they offer valuable insights for understanding underlying motivations and guiding further security analysis.
  • Graph-level Understanding. LLMs can effectively extract node attributes and synthesize meaningful overviews of transaction graphs, enabling the identification of complex patterns and relational structures. Compared with traditional algorithms, which often struggle with high-dimensional and context-rich data, LLMs demonstrate a stronger capacity to process raw graph representations and capture intricate dependencies and anomalies.
  • Support for Cybercrime and Security Analysis. LLMs can assist in identifying anomalous transaction patterns, inferring the potential motivations behind suspicious activities, and providing interpretable reasoning processes. These capabilities are particularly valuable in security-critical contexts such as anti-money laundering, fraud detection, and darknet market tracing, offering law enforcement and researchers additional clues and explainable evidence.
Remaining Challenges. We also outline three key challenges in applying LLMs to cryptocurrency transaction graphs.
  • Token Limits. They restrict the amount of graph data that can be processed at once, hindering the analysis of large cryptocurrency transaction graphs. With only limited portions available, LLMs often lack sufficient context to deliver accurate insights, making it difficult to capture the complexity and nuances of the data.
  • Reference Graph Selection. Choosing representative labeled samples is challenging, as different subgraphs emphasize different features and may introduce bias. Such choices directly affect classification outcomes and their explanations. The problem is further compounded by token limits, which restrict the number of reference samples that can be included, especially for raw graph data, thereby constraining the model’s ability to generalize.
  • Explanation Accuracy. Improving the reliability of LLM-generated explanations remains a major challenge. Current outputs may contain incomplete reasoning or subtle inaccuracies, which limits their utility for high-stakes applications such as illicit activity detection. Developing more rigorous prompting strategies or integration with external knowledge sources could enhance the fidelity of these explanations. Achieving this would not only improve interpretability but also strengthen the overall effectiveness of cryptocurrency transaction analysis.
VI-C Factors Influencing LLM Performance
The performance of LLMs on transaction graphs is shaped by multiple factors, with the first being the model bottleneck. Although major updates and optimizations can enhance the capability of LLMs in certain aspects, they may not resolve the inherent limitations of LLMs in transaction graph understanding. For example, when using the same graph feature data for classification, GPT-4 and GPT-4o show only slight improvements over GPT-3.5. At present, a more effective way to boost performance may involve increasing the number of labeled samples and enriching the feature set in graph feature-based classifications, which also require fewer tokens compared to processing raw graphs.
Another factor lies in the size and complexity of transaction graphs. Smaller graphs tend to yield better results on foundational metrics, while larger graphs may reduce effectiveness. At the contextual interpretation level, we observed that even when using the same data, accuracy may vary slightly, by around 5%. Moreover, due to the application of LLM4TG, there is a slight loss of temporal information, which could affect results when raw graphs are used.
Finally, the representation of data and the choice of features also play a crucial role. The type of input, such as graph features versus raw graphs, and the availability of labeled samples directly influence performance. Graph feature-based data is generally more efficient in terms of LLM token consumption, and its classification effectiveness can be further improved through the inclusion of additional features and labels.
VII Conclusion
In addressing our research questions, we find that (RQ1) the LLM4TG format, combined with CETraS, reduces token redundancy and preserves graph structure, enabling efficient analysis of large Bitcoin graphs; (RQ2) our three-level framework effectively measures LLMs’ capacity to understand transaction graphs, capturing both local node details and broader behavioral patterns; and (RQ3) while engineered features simplify analysis through pre-computed metrics, raw graph data allows deeper insights and higher potential accuracy, but also exposes larger performance disparities across models.
Acknowledgment
This research was supported by the OpenAI Researcher Access Program (Project ID: 0000007730).